Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 29,000 39,000 68,000
Fixed manufacturing overhead costs $ 730,000 $ 270,000 $ 1,000,000
Variable manufacturing overhead cost per machine-hour $ 6.10 $ 6.10

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 373,000 $ 323,000 $ 696,000
Direct labor cost $ 230,000 $ 160,000 $ 390,000
Machine-hours 19,000 10,000 29,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 230,000 $ 300,000 $ 530,000
Direct labor cost $ 120,000 $ 270,000 $ 390,000
Machine-hours 10,000 29,000 39,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

2. Assume Delph uses departmental predetermined overhead rates based on machine-hours.

a. Compute the departmental predetermined overhead rates.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 120% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?
Answer #1
a
Predetermined
overhead rate
Molding Department 31.27 =6.1+(730000/29000)
Fabrication Department 13.02 =6.1+(270000/39000)
b
Job D-70 Job C-200
Direct materials cost 696000 530000
Direct labor cost 390000 390000
Overhead:
Molding Department 594130 312700
Fabrication Department 130200 377580
Total manufacturing cost 1810330 1610280
c
Job D-70 Job C-200
Total manufacturing cost 1810330 1610280
Add: Markup @20% 362066 322056
Bid prices 2172396 1932336
d
Total manufacturing cost Job D-70 1810330
Total manufacturing cost Job C-200 1610280
Cost of goods sold 3420610
Workings:
Overhead applied:
Job D-70
Molding Department = 19000*31.27 = 594130
Fabrication Department = 10000*13.02= 130200
Job C-200
Molding Department = 10000*31.27 = 312700
Fabrication Department = 29000*13.02=377580